How many atoms of oxygen are there in 45.67 g of C6H12O6?
1
Expert's answer
2020-02-17T07:10:09-0500

Solution.


Calculate molecular mass C6H12O6:

"M=12*6+12+16*6=180" g/mole.


Calculate amount mole of substance in 45.67 g of C6H12O6:

"\nu=\dfrac{m}{M}=\dfrac{45.67}{180}=0.25" mole.


1 mole contains number of molecules "N_A=6.02*10^{23}" molecules, so in 45.67 g C6H12O6 contains number of molecules:

"N=N_A*\nu=6.02*10^{23}*0.25=1.505*10^{23}"molecules.


1 molecule contains 6 atoms of oxygen, so number of atoms:

"N_O=N*6=1.505*10^{23}*6=9.03*10^{23}" atoms of oxygen.


Answer: "9.03*10^{23}" atoms of oxygen.